The perimeter of a rectangular garden is 98 feet. The length of the garden is 4 feet more than twice the width. What is the length of the garden.
Answer by macston(5194)

W=width; L=length=2W+4ft P=perimeter=2(L+W)=98ft
P=2(L+W)
98ft=2(L+W) Divide each side by 2.
49ft=L+W Substitute for L.
49ft=2W+4ft+W Subtract 4 ft from each side.
45ft=3W Divide each side by 3.
15ft=W
L=2W+4ft=2(15ft)+4ft=34ft ANSWER: The length is 34 feet.
CHECK:
P=2(L+W)
98ft=2(34ft+15ft)
98ft=2(49ft)
98ft=98ft